# Destabilisation of local magnetic anisotropy in heavy-fermion compounds

### v2

@@ Abstract
Our results are shown in Fig. 3 for two representative fillings. We can see that order takes place along the high-temperature hard-axis at strong enough Kondo coupling, when the excitation energy for interband particle-hole fluctuations is sufficiently large and intraband fluctuations dominate. $\chi_{cf}(\mathbf{q})$ is manifestly positive in this limit (and $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}$ at the ordering wavevector if $\chi_{cf}(\mathbf{q})$ and $\chi_{cc}(\mathbf{q})$ are maximised at similar wavevectors). Stronger Kondo coupling (at fixed $N\gg 1$ in our theory) corresponds to a greater dominance of the Kondo energy scale over the RKKY energy scale and thus more itinerant magnetism. We can indeed see that compounds with higher ratios of $T^{*}_{\rm coh}/T_{N,C}$ are more likely to exhibit order along the high-temperature hard-axis.
At weak coupling interband particle-hole fluctuations dominate. They enforce the antiferromagnetic correlation between conduction and $f_{1}$-electrons, and $\chi_{cf}(\mathbf{q})<0$ (and $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}<0$ at the ordering wavevector if $\chi_{cf}(\mathbf{q})$ and $\chi_{cc}(\mathbf{q})$ are extremised at similar wavevectors). These fluctuations are the only allowed ones at $T=0$ in a Kondo insulator, where applying a small field to conduction electrons cannot change the occupancy of up or down heavy-fermions and can only make the singlet state slightly asymmetric, which leads to opposing magnetisations of $c$ and $f_{1}$ moments. Because the indirect band gap is smaller than the direct one, the antiferromagnetic $\mathbf{q}=(\pi/a,\pi/a,\pi/a)$ phase also dominates in this limit. (We have observed ferromagnetic order along the high-temperature easy axis only at low fillings.) The fact that the coupling needs to be strong enough (or filling low enough) to observe ferromagnetism probably explains why it is less common.
On the other hand, when intraband particle-hole fluctuations close to the Fermi surface dominate, the dominant effect from an external field applied only to conduction electrons is to increase (decrease) the occupancy of up (down) heavy-fermions, and hence, because $f_{1}$ and $c$ moments share the same $z$-angular momentum in the heavy-fermion quasiparticle they acquire the same magnetisation and $\chi_{cf}(\mathbf{q})>0$. At strong coupling, Kondo hybridisation, being angular momentum conserving, favours magnetic alignment (in applied field) of $f_{1}$ and $c$ moments, whereas the RKKY interaction favours antialignment of $f_{2}$ and $c$ moments. A weak single-ion anisotropy, which favours antialignment of $f_{1}$ and $f_{2}$ moments along the high-temperature hard axis, thus lowers (increases) the energy of a state that magnetically orders along the high-temperature hard (easy) axis by stabilising the coexistence of Kondo hybridisation and RKKY interactions. While ferromagnetic and antiferromagnetic magnetic order both can take place along the high-temperature hard axis at sufficiently strong coupling, where intraband particle-hole fluctuations close to the Fermi surface dominate, the latter is favoured when a finite momentum transfer connects Fermi surface points of the same energy, and is thus lattice and filling specific by nature, and perhaps therefore less common.
@@
Conclusion – In this Letter, we have shown that the development of flat bands in heavy-fermion metals can destabilise the single-ion magnetic anisotropy of the local moments, that sets the easy axis at high temperatures, by generating an opposing exchange anisotropy. The destabilisation leads to a reorientation of the easy-axis with cooling and magnetic order (if it takes place) at 90 degrees to the high-temperature easy-axis. The destabilisation is mostly likely to occur in more itinerant heavy-fermion compounds that have a clear resistivity maximum at a temperature significantly higher than the magnetic ordering temperature. Our theoretical analysis follows an extension of the Read-Newns theory to the underscreened Kondo lattice, which naturally gives rise to the coexistence of local moments and Kondo hybridisation, even when the energy scale of the latter is much greater. In agreement with our theory, a wide range of heavy-fermion compounds exhibit a crossing of uniform susceptibilities at a temperature that tracks the coherence energy scale, a trend not previously noticed.
@@
where $\bar{S}^{\eta}_{2\alpha i}$ have been promoted to spin-$1/2$ operators, constant terms have been dropped, $\chi_{ff}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{S}^{\eta}_{1\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{S}_{1\alpha}^{\eta}(\mathbf{0})\rangle_{H_{0}}$, $\chi_{cf}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{S}^{\eta}_{1\alpha}(\mathbf{0})\rangle_{H_{0}}$, $\chi_{cc}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{0})\rangle_{H_{0}}$, and $\chi(\mathbf{q})=\sum_{i}\chi(\mathbf{r})e^{-i\mathbf{q}\cdot\mathbf{r}_{i}}$. These susceptibilities can be easily evaluated to obtain the expressions given in the main text once $H_{0}$ is expressed in terms of the heavy-fermion quasiparticle creation/annihilation operators: $\sum_{\nu\mathbf{k}\alpha\sigma}E^{\nu}_{\mathbf{k}}\left(u_{\mathbf{k}}^{\nu}c_{\sigma\alpha}(\mathbf{k})+v_{\mathbf{k}}^{\nu}f_{1\sigma\alpha}(\mathbf{k})\right)^{\dagger}\left(u_{\mathbf{k}}^{\nu}c_{\sigma\alpha}(\mathbf{k})+v_{\mathbf{k}}^{\nu}f_{1\sigma\alpha}(\mathbf{k})\right)=\sum_{\nu\mathbf{k}\alpha\sigma}E^{\nu}_{\mathbf{k}}d^{\dagger}_{\nu\alpha\sigma}(\mathbf{k})d_{\nu\alpha\sigma}(\mathbf{k})$, where
@@
Appendix B Susceptibility crossings in compounds with orthorhombic symmetry
The effective Hamiltonian in Eq. 7 would suggest that the axis with the lowest high-temperature susceptibility (largest $D^{\eta}$) should always become the low-temperature easy axis. This is seen in tetragonal systems (see Table I) but rarely in orthorhombic systems (except YbRhSb Umeo et al. (2012) at high pressures). In most compounds, only the two highest susceptibilities cross with the third remaining quenched. We suggest that, in those cases, the anisotropy of the axis with the smallest susceptibility is too large for the Kondo hybridisation to overcome. In fact, it usually has the lowest effective moment as well as the lowest Curie-Weiss constant, which suggests that the mechanism is not in play. $H_{\rm eff}$ is only valid for $D^{\eta}\ll T_{K}$, but if we increase the anisotropy and look at the optimised variational wavefunction for the ordered ground state, the hybridisation field that is responsible for the reduction of the effective moment and the reinforcement of the Curie-Weiss constant eventually collapses and we revert to free-moment behaviour, where the higher the $D^{\eta}$, the lower the susceptibility. This happens before the hybridisation field collapses for the two directions with lower $D^{\eta}$. We note that, although the collapse signals the return of the free-moment state, it is beyond the validity of our perturbative model.

### v3

@@ Abstract
Our results are shown in Fig. 3 for two representative fillings. We can see that order takes place along the high-temperature hard-axis at strong enough Kondo coupling, when the excitation energy for interband particle-hole fluctuations is sufficiently large and intraband fluctuations dominate. $\chi_{cf}(\mathbf{q})$ is manifestly positive in this limit (and $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}$ at the ordering wavevector if $\chi_{cf}(\mathbf{q})$ and $\chi_{cc}(\mathbf{q})$ are maximised at similar wavevectors). Stronger Kondo coupling (at fixed $N\gg 1$ in our theory) corresponds to a greater dominance of the Kondo energy scale over the RKKY energy scale and thus more itinerant magnetism. We can indeed see that compounds with higher ratios of $T^{*}_{\rm coh}/T_{N,C}$ are more likely to exhibit order along the high-temperature hard-axis. While ferromagnetic and antiferromagnetic magnetic order both can take place, the latter is favoured when a finite momentum transfer connects Fermi surface points of the same energy, and is thus lattice and filling specific by nature, and perhaps therefore less common.
At weak coupling interband particle-hole fluctuations dominate. They enforce the antiferromagnetic correlation between conduction and $f_{1}$-electrons, and $\chi_{cf}(\mathbf{q})<0$ (and $\chi_{cf}(\mathbf{q})-\bar{\chi}_{cf}<0$ at the ordering wavevector if $\chi_{cf}(\mathbf{q})$ and $\chi_{cc}(\mathbf{q})$ are extremised at similar wavevectors). Because the indirect band gap is smaller than the direct one, the antiferromagnetic $\mathbf{q}=(\pi/a,\pi/a,\pi/a)$ phase also dominates in this limit.
@@
Conclusion – In this Letter, we have shown that the development of flat bands in heavy-fermion metals can destabilise the single-ion magnetic anisotropy of the local moments, that sets the easy axis at high temperatures, by generating an opposing exchange anisotropy. The destabilisation leads to a reorientation of the easy-axis with cooling and magnetic order (if it takes place) at 90 degrees to the high-temperature easy-axis. The destabilisation is mostly likely to occur in more itinerant heavy-fermion compounds that have a clear resistivity maximum at a temperature significantly higher than the magnetic ordering temperature.
@@
where $\bar{S}^{\eta}_{2\alpha i}$ have been promoted to spin-1/21/21 / 2 operators, constant terms have been dropped, $\chi_{ff}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{S}^{\eta}_{1\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{S}_{1\alpha}^{\eta}(\mathbf{0})\rangle_{H_{0}}$, $\chi_{cf}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{S}^{\eta}_{1\alpha}(\mathbf{0})\rangle_{H_{0}}$, $\chi_{cc}(\mathbf{r})=\int_{0}^{\beta}\langle e^{H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{r})e^{-H_{0}\tau}\hat{s}^{\eta}_{\alpha}(\mathbf{0})\rangle_{H_{0}}$, and $\chi(\mathbf{q})=\sum_{i}\chi(\mathbf{r})e^{-i\mathbf{q}\cdot\mathbf{r}_{i}}$. These susceptibilities can be easily evaluated to obtain the expressions given in the main text once $H_{0}$ is expressed in terms of the heavy-fermion quasiparticle creation/annihilation operators: $\sum_{\nu\mathbf{k}\alpha\sigma}E^{\nu}_{\mathbf{k}}\left(u_{\mathbf{k}}^{\nu}c_{\sigma\alpha}(\mathbf{k})+v_{\mathbf{k}}^{\nu}f_{1\sigma\alpha}(\mathbf{k})\right)^{\dagger}\left(u_{\mathbf{k}}^{\nu}c_{\sigma\alpha}(\mathbf{k})+v_{\mathbf{k}}^{\nu}f_{1\sigma\alpha}(\mathbf{k})\right)=\sum_{\nu\mathbf{k}\alpha\sigma}E^{\nu}_{\mathbf{k}}d^{\dagger}_{\nu\alpha\sigma}(\mathbf{k})d_{\nu\alpha\sigma}(\mathbf{k})$, where
@@
Appendix B Metamagnetism
We mention here the apparent correlation between metamagnetism and the crossing of uniform susceptibilities. As shown in Table I, compounds that exhibit metamagnetism along the high-temperature easy axis (HTEA) also exhibit a crossing of the uniform susceptibilities at a temperature $T_{\rm cross}$, as well as a maximum along the HTEA at a temperature $T^{\rm hard}_{\rm max}<T_{\rm cross}$.